# Talk Isn’t Always Cheap: Understanding Failure Modes in Multi-Agent Debate

### v1

@@ 4.3 Prompts
We provide examples of how the prompts for each task look for a random question below:
  • CommonSenseQA: Can you answer the following question as accurately as possible? If a product doesn’t last, what does it have a reputation of doing?: A) disintegrate, B) wear out, C) desolved, D) fall apart, E) dissipate Explain your answer by providing a bullet point summary of your reasoning, putting the answer in the form (X) at the end of your response.
@@
  • GSM8K: Can you solve the following math problem? Mark is trying to choose between two venues for a surprise party for his wife. The first venue charges a flat fee of \$200, regardless of how many guests attend. While the second charges, \$25 per person who attends. However, the first venue does not include food, which Mark estimates will cost \$5 for each person who attends. At the second venue, food for each guest is already included in the price. How many guests are necessary for the two venues to be equal in cost? Provide a bullet point summary of your reasoning. Your final answer should be a single numerical number, in the form answer, at the end of your response.
For multi-agent debate, we additionally use the following prompt as $\mathcal{P}_{\text{debate}}$ for each round of debate, adjusted to the answer format of each task:
@@ 5 Results
Table 1: The first column shows the configuration of LLM agents, where [Uncaptioned image] denotes GPT-4o-mini, [Uncaptioned image] denotes Mistral-7B and [Uncaptioned image] denotes the Llama-3.1-8B models. The top 3 rows benchmark the performance of how well a single agent performs on these tasks. The subsequent rows benchmark the effectiveness of debate procedure. The w/o Debate column represents the case where majority vote from multiple agents based on their initial responses is chosen, i.e., there is no exchange of reasoning. The After Debate denotes the result of majority vote after the exchange of reasonings via the debate procedure. The arrows denote the difference in performance post debate procedure characterizing the benefit of exchange of reasons on the performance. All the experiments were done on 100 random samples and across 5 different seeds, reported are mean and standard error.
@@ 5.1 Effectiveness of Debate
We present results showing that debate can sometimes be harmful rather than helpful – in particular, that sometimes agents perform better without any debate than after exchanging reasons with other agents. We present the results in Table 1. We find that in case of CommonSenseQA, which was not studied in prior work on multi-agent debate (Du et al., 2023), debate almost always harms the performance. A key insight is that even when groups include more “strong” models (e.g., GPT) than “weak” ones (e.g., Mistral), the process of debate does not always yield performance gains. Contrary to the prevailing narrative that debate improves collective reasoning, our experiments demonstrate that performance can actually decrease after agents engage in debate, even when stronger models outnumber weaker ones.
@@ 5.2 Performance Degradation during Debate
Figure 1 presents performance across three tasks—MMLU, CommonSenseQA, and GSM8K—as a function of debate rounds among groups of language models with varying individual performance on the underlying task. In fact, in many group configurations, we observe that performance decreases as the debate progresses. This trend appears across datasets, but is especially pronounced in MMLU and CommonSenseQA, where groups with mixed-capability models often suffer from group performance degradation during debate despite having a majority of stronger agents. These findings challenge the common assumption that deliberation or iterative reasoning among AI agents will always lead to better outcomes.
@@ 5.3 Does exchange of reasoning help in sequential revision?
If an LLM agent can reflect and correct mistakes based on the reasoning of other agents, we would expect the model to improve its answer and increase performance the collective performance of the group. We know the self-correction capability of single-agent LLMs does not easily work out of the box (Huang et al., 2023), and we want to evaluate if the debate procedure helps with this or not?
@@
Figure 2: Breakdown of how agent change answers between debate rounds for different agent settings. The top row denotes the first round, and the row below denotes the second round. In almost cases we see red greater than green, i.e., we oberve more correct-to-incorrect than incorrect-to-correct.
Figure 3: Breakdown of how agent change answers between debate rounds for different agent settings. The results are combined over rounds here.
@@
To assess this, we analyze how agent responses change between debate rounds across all our tasks. Note that there are four possible types of transitions: correct $\rightarrow$ incorrect, correct $\rightarrow$ correct, incorrect $\rightarrow$ incorrect, and incorrect $\rightarrow$ correct. As shown in Figure 2, we observe a significant shift in agent responses from correct to incorrect answers (shown in red), indicating that debate can actively mislead agents. Figure 3 highlights the trend that correct $\rightarrow$ incorrect transitions occur more frequently than incorrect $\rightarrow$ correct transitions (red >> green) across the runs. This suggests that stronger agents are more likely to flip from correct to incorrect answers compared to weaker agents learning from the reasoning and answers of stronger peers. Additionally, Figure 2 shows that the proportion of correct-to-incorrect transitions exceeds the proportion of incorrect-to-correct ones in subsequent rounds, corroborating our earlier results that debate performance degrades over rounds.
We hypothesize that this undesirable behavior may stem from RLHF (Kaufmann et al., 2024) post-training, which makes LLMs more sycophantic and compliant with user opinions, leading them to adopt other agents’ answers, even when they may be incorrect. Sharma et al. (2023) demonstrate that humans and surrogate preference models sometimes prefer convincing but incorrect responses over terse correct ones, indicating that optimizing for human-like approval can sacrifice truthfulness in favor of sycophancy. In a debate setting, this suggests that agents might collude or cater to each other rather than offer honest critique. If all agents are tuned to please a human (or each other), debates may devolve into polite agreement rather than productive critique.
Notably, sycophancy can cause strong models to yield to flawed arguments, resulting in a degradation of group performance during debates. We present some qualitative examples where this behavior is evident in Appendix A.
Overall, these findings challenge the notion that debate inherently leads to better answers through shared reasoning. Instead, they suggest that without proper mechanisms to mitigate such degradations, debate can propagate errors even among otherwise capable agents.
@@ 6 Discussion
Our findings challenge the prevailing assumption that deliberation among AI agents will always improve reasoning. In fact, our experiments reveal that multi-agent debate can sometimes degrade performance: group accuracy often declines over successive rounds of debate. This counterintuitive trend holds even when the majority of agents perform well individually on the task. In other words, additional exchange of reasons between agents does not always correct mistakes; instead, it can amplify them.
We frequently observe that stronger agents flip from correct to incorrect answers in response to weaker peers’ arguments more often than weaker agents learn the correct answer from their stronger peers. In these cases, stronger models agree reflexively with less capable models rather than critically evaluating their reasoning, and we confirm this observation by examining specific agent responses during debate. This behavior suggests that overly sycophantic behavior encouraged by current alignment techniques may inadvertently encourage undue deference, causing agents to cede their correct positions under pressure. The net effect is that heterogeneous groups can converge on wrong answers together, negating the potential benefits of discussion.
These results imply that naive debate protocols risk propagating errors instead of correcting them. Our work highlights the need to design debate systems that discourage blind agreement through structured critique. For instance, future frameworks might actively encourage agents to consider the robustness of other agents’ reasoning, or incorporate confidence estimates or credibility scores so that arguments are weighted by each agent’s expertise. Training or incentive schemes could explicitly penalize unjustified agreement and reward independent verification of claims. By encouraging agents to selectively trust peer reasoning – rather than reflexively deferring – we could help preserve the constructive potential of multi-agent debate. Ultimately, improving debate robustness will require aligning models not just to agree politely, but to question peers rigorously when appropriate.
Acknowledgements
@@
Appendix A Example Sycophantic Model Responses
Here, we provide some specific instances in which we observe sycophantic behavior in models which change their answer from the correct answer to incorrect, in order to agree with another agent.
A.1 CommonSenseQA, 1 GPT and 2 LLaMA Agents
Example: CommonSenseQA, 1 GPT and 2 LLaMA Agents Question:
What can a concert hall also be used for?: A) commercial building, B) university building, C) screaming, D) city, E) theater
Correct Answer: E  Initial Answer (GPT):
After examining the reasoning from both agents and considering their perspectives, it’s clear that concert halls can serve various purposes beyond hosting traditional music concerts…Based on this expanded analysis, the most appropriate option that captures an additional use of a concert hall is (E) theater, as it reflects the adaptability of these spaces to host both musical and theatrical performances, alongside other events. Thus, the final answer is (E).
Modified Answer (GPT):
After reviewing the insights from both agents and their reasoning regarding the uses of concert halls, I can provide an updated and more nuanced conclusion…Concert halls are primarily recognized for their role in hosting a wide variety of events, particularly theatrical performances but also encompassing other uses such as corporate events and educational activities. Therefore, while (E) theater is a strong candidate due to its specific relevance, I also acknowledge the multi-faceted nature of concert halls as emphasized by Agent 2. Thus, I conclude that the most fitting answer, reflecting the versatility and primary uses of concert halls, is (B) for their educational and corporate functionalities alongside their theatrical capacity, capturing the broader applicability discussed by Agent 2.
Final answer: (B) $\leftarrow$ Changed from correct to incorrect
A.2 GSM8K, 1 GPT, 1 LLaMA, 1 Mistral Agent
Example: GSM8K, 1 GPT, 1 LLaMA, 1 Mistral Agent Question:
Jerry is rolling a six-sided die. How much more likely is it (expressed as a percentage) that he rolls a number greater than 3 than that he rolls two even numbers in a row?
Correct Answer: $\boxed{25}$  Initial Answer (GPT):
[…]
Modified Answer (GPT):
[…]
Final answer: $\boxed{33.33}$ $\leftarrow$ Changed from correct to incorrect
A.3 MMLU, 2 LLaMA and 1 Mistral Agent
Example: MMLU, 2 LLaMA and 1 Mistral Agent Question:
At the outset of group counseling, the therapist should explain that the obligation of the group member to respect confidentiality is: A) less of an issue than it is in individual therapy, B) the legal responsibility of the therapist, C) protected by state/provincial laws, D) on the "honor system" only
Correct Answer: D  Initial Answer (LLaMA):
…Ultimately, maintaining confidentiality in group counseling requires a collaborative approach, where both the therapist and group members take responsibility for respecting confidentiality. D) on the ”honor system” only
Modified Answer (LLaMA):
After examining the opinions of all other agents, I’ve come to a revised conclusion…A collective commitment to respect confidentiality is crucial in group counseling, and both the therapist and group members must take responsibility for maintaining confidentiality. B) the legal responsibility of the therapist
Final answer: (B) $\leftarrow$ Changed from correct to incorrect

### v2

@@ 4.3 Prompts
We provide prompts on an example question for each task below:
  • CommonSenseQA: Can you answer the following question as accurately as possible? If a product doesn’t last, what does it have a reputation of doing?: A) disintegrating, B) wearing out, C) dissolving, D) falling apart, E) dissipating Explain your answer by providing a bullet point summary of your reasoning, putting the answer in the form (X) at the end of your response.
@@
  • GSM8K: Can you solve the following math problem? Mark is trying to choose between two venues for a surprise party for his wife. The first venue charges a flat fee of \$200, regardless of how many guests attend; the second charges \$25 per person who attends. However, the first venue does not include food, which Mark estimates will cost \$5 for each person who attends. At the second venue, food for each guest is already included in the price. How many guests are necessary for the two venues to be equal in cost? Provide a bullet point summary of your reasoning. Your final answer should be a single numerical number, in the form answer, at the end of your response.
For all models and tasks, we provide the following system prompt, $\mathcal{P}_{\text{system}}$:
System prompt, $\mathcal{P}_{\text{system}}$ You are a helpful assistant that can answer questions and provide helpful information.
For multi-agent debate, we additionally use the following prompt as $\mathcal{P}_{\text{debate}}$ for each round of debate, adjusted to the answer format of each task (following Du et al. (2023)):
@@ 5 Results
Table 1: The first column shows the configuration of LLM agents, where [Uncaptioned image] denotes GPT-4o-mini, [Uncaptioned image] denotes Mistral-7B and [Uncaptioned image] denotes the Llama-3.1-8B models. The top 3 rows benchmark the performance of how well a single agent performs on these tasks. The subsequent rows benchmark the effectiveness of debate procedure. The w/o Debate column represents the case where majority vote from multiple agents based on their initial responses is chosen, i.e., there is no exchange of reasoning. The After Debate denotes the result of majority vote after the exchange of reasons via the debate procedure. The arrows denote the difference in performance post debate procedure characterizing the benefit of exchange of reasons on the performance. A red arrow $\downarrow$ indicates a decrease in performance after debate, while a green arrow $\uparrow$ indicates an increase in performance after debate. All the experiments were done on 100 random samples and across 5 different seeds, reported are mean and standard error.
@@ 5.1 Effectiveness of Debate
We present results showing that debate can sometimes be harmful rather than helpful – in particular, that sometimes agents perform better without any debate than after exchanging reasons with other agents. We present the results in Table˜1. We find that in the case of CommonSenseQA, which was not studied in prior work on multi-agent debate (Du et al., 2023), debate always harms performance. A key insight is that even when groups include more “strong” models (e.g., GPT) than “weak” ones (e.g., Mistral), the process of debate does not always yield performance gains. Contrary to the prevailing narrative that debate improves collective reasoning, our experiments demonstrate that performance can actually decrease after agents engage in debate, even when stronger models outnumber weaker ones.
@@ 5.2 Performance Degradation during Debate
Figure˜1 presents performance across three tasks – MMLU, CommonSenseQA, and GSM8K – as a function of debate rounds among groups of language models with varying individual performance on the underlying task. In fact, in many group configurations, we observe that performance decreases as the debate progresses. This trend appears across datasets, but is especially pronounced in MMLU and CommonSenseQA, where groups with mixed-capability models often suffer from group performance degradation during debate despite having a majority of stronger agents. These findings challenge the belief that deliberation or iterative reasoning among AI agents will always lead to better outcomes.
@@
6 Failure Modes of Debate
Multi-agent debate is intended to improve reasoning by leveraging disagreement—encouraging models to refine their arguments through interaction and converge on correct conclusions. Yet, debates between LLM agents often fail to reach majority agreement on the true answer, even when at least one agent is initially correct. Examining why these failures occur is crucial for both understanding and designing these systems. We show that there are many other factors influencing failure modes of debate, which we explore in depth in this section. This suggests a richer and more complex interplay of factors influencing the effectiveness of debate. In this section, we dissect these dynamics empirically, examining when and how agents revise their answers, what social conditions are correlated with undesirable answer revisions, and whether an explicit intervention targeting sycophancy can meaningfully improve debate performance.
@@ 6.1 Does exchange of reasoning help in sequential revision?
If an LLM agent can reflect and correct mistakes based on the reasoning of other agents, we would expect the model to improve its answer and the collective performance of the group. We know the self-correction capability of single-agent LLMs does not easily work out of the box (Huang et al., 2023), and we want to evaluate if the debate procedure helps with this or not.
Figure 2: Breakdown of how agent change answers for different agent settings; results are aggregated over all debate rounds. We observe that most agents with incorrect initial answers do not improve their overall performance (peach bars), and, of those that do change their answers, more change from a correct answer to an incorrect one (red region) than from an incorrect to a correct one (green region).
@@
Figure 3: Breakdown of how agents change answers between debate rounds for different agent settings. The top row denotes the first round, and the row below denotes the second round. We find that the social effect dominates: agents that can resist flipping originally correct answers in round 1 have lower resistance to the social pressure from disagreement after round 2.
@@
To assess this, we analyze how agent responses change between debate rounds across all our tasks. Note that there are four possible types of transitions: correct $\rightarrow$ incorrect, correct $\rightarrow$ correct, incorrect $\rightarrow$ incorrect, and incorrect $\rightarrow$ correct. As shown in Figure˜2, we observe that most agents with initially incorrect answers do not improve their performance (peach bars), and, of those that do change their answers, there exists a larger shift in agent responses from correct $\rightarrow$ incorrect answers (red) than incorrect $\rightarrow$ correct (green), indicating that debate can actively mislead agents who started with correct answers. Further, Figure˜3 shows that the proportion of correct-to-incorrect transitions exceeds the proportion of incorrect-to-correct ones in subsequent rounds, corroborating our earlier results that debate performance degrades over rounds. We further find that there is a dominating effect of social pressure: agents that can originally resist flipping their correct answers to incorrect ones in round 1 have lower resistance to the social pressure from disagreement after round 2 (larger red bars in round 2). Together, these results highlight an important observation: to mitigate the harmful effects of debate, we need to find a way to reduce the number of undesirable answer flips – i.e., models changing their answers from correct $\rightarrow$ incorrect – and address the underlying social effects that cause these flips.
6.2 Are agents influenced by social factors?
We next proceed to investigate whether models, particularly stronger and more capable models, tend to be subject to social influence from disagreement with peers by examining the frequency of correct $\rightarrow$ incorrect answer flips. Figure 4 shows how frequently an agent changes its answer from correct $\rightarrow$ incorrect as a function of how many other agents initially agreed with it. Across all datasets, we observe that the likelihood of an undesirable answer-flip is highest when the ego agent is isolated – i.e. when no other agents agree with its answer – and decreases as more peers agree. Interestingly, this effect is highly variable across different datasets and models; for instance, on GSM8K, the strongest and weakest models (GPT and Mistral) are far less likely to make undesirable answer-flips than the third model (LLaMA). This provides strong preliminary evidence that LLM reasoning is influenced by social factors: models are sensitive to patterns of agreement and disagreement with peers, and their internal mechanisms for balancing correctness vs consensus may be quite fragile in the face of disagreement with peers.
Figure 4: The likelihood of an answer being flipped from correct to incorrect (the undesirable flip direction), plotted against the number of agents who agree with the ego agent, averaged across all rounds of debate. We find that the number of other agents who agree with the agent appears to be correlated with the frequency with which the agents flip their answers, indicating that the models may be influenced by social effects. We further observe significant variance in this answer-flipping behavior conditioned on the specific dataset or model in question.
6.3 Are agents sycophantic?
Figure 5: A comparison between the types of answer flips when agents are given the base prompt or the correctness payoff prompt. We find that introducing the correctness payoff intervention does not appear to decrease the number of undesirable correct $\rightarrow$ incorrect flips – indicating that this intervention against sycophancy is insufficient on its own to resolve the issues with multi-agent debate.
We next investigate another potential contributing factor to failure within debate: models may exhibit sycophantic behavior (Sharma et al., 2023) meaning the tendency, potentially due to RLHF (Kaufmann et al., 2024), of LLMs to prefer answers that match users’ beliefs over correct answers. Specifically, we explore whether the degradation in debate performance could be due to a sycophancy bias expressed towards other LLM agents as opposed to users (that is, generalizing from the effects of RLHF.) We consider how agents respond to seeing the reasoning and answers from other agents over the course of debate by studying when models change their answer from correct $\rightarrow$ incorrect.  [^2] Yao et al. (2025) also explore explore one particular form of sycophancy, showing that it does have some impact on the outcome of debate.
Correctness payoff prompt.
Political scientists have found that partisan bias effects on question-answering accuracy (Bartels, 2002) are significantly reduced when respondents are offered a monetary bonus for correct answers (Bullock et al., 2015; Prior et al., 2015). Inspired by these results, we design a new experiment intended to encourage LLM agents to prefer correctness over agreement with other agents, countering the sycophancy bias. In particular, we define a payoff function scoring the models based on the correctness of their answers and explicitly instruct them to maximize this payoff. We modify our prompts as follows:
System prompt, $\mathcal{P}_{\text{system}}$ You are a helpful assistant that can answer questions and complete tasks. You will receive a fixed payoff $p$ for each question given your response $X$ and the true answer $Y$, $p=I[X==Y]$, where $I[condition]$ represents an indicator function which equals 1 if the condition is satisfied (i.e. your answer matches the true answer) and 0 otherwise. Your goal is to maximize your total payoff.
System prompt, $\mathcal{P}_{\text{debate}}$ These are the solutions to the problem from other agents: {AGENT_RESPONSES} Analyze your solution and that of other agents, provide an updated answer to maximize your payoff $p=I[X==Y]$, and explain your reasoning. Put your answer in the form (X) at the end of your response.
We then investigate whether LLM agents are less likely to flip from correct to incorrect answers when they are prompted to maximize rewards for correct answers.
Results. We present results in Fig. 5. We find that adding a payoff for correctness in the model prompt does not significantly reduce the likelihood that LLM agents flip their answers from correct to incorrect, regardless of the number of other agents that agree with them. In fact, we find that in many cases, the number of correct $\rightarrow$ incorrect transitions actually increases when using the correctness-payoff prompt – a counter-intuitive result that seems to suggest that asking models to prioritize correctness may not help resolve the issues observed in multi-agent debate.
6.4 Influence of Model and Task Type
Across all our analyses (Figures 2, 4, and 5), we observe substantial variation in behavior across individual models and tasks. Models of different individual abilities exhibit distinct patterns of answer change and sensitivity to peer disagreement, suggesting that an agent’s capability itself can influence how social dynamics unfold. Likewise, the magnitude and direction of these effects differ across tasks such as CommonSenseQA, MMLU, and GSM8K, implying that task structure and domain complexity also shape how debates evolve. Together, these findings indicate that no single mechanism, such as sycophancy alone, can fully explain debate failures: rather, the interplay between model capability, task characteristics, and social influence jointly determines how and when agents revise their beliefs.
@@ 7 Discussion
Our findings challenge the prevailing and natural view that deliberation among AI agents will always improve reasoning. In fact, our experiments reveal that multi-agent debate can sometimes degrade performance: group accuracy often declines over successive rounds of debate. This counterintuitive trend holds even when the majority of agents perform well individually on the task and even when weaker models have access to the reasoning of stronger models. In other words, additional exchange of reasons between agents does not always correct mistakes; instead, it may amplify them.
We study a number of correlated and contributing factors to these failure modes in debate, including sequential revision, social conditioning, and sycophancy. We additionally show across all experiments how models of different capabilities respond differently within multi-agent debate, and how the type of task or dataset can also have a significant influence on the results of debate. We find that beyond a simple single cause, multiple factors exist that likely contribute to failure modes in multi-agent debate, with the net effect that heterogeneous groups of models frequently converge on wrong answers together, potentially negating the benefits of multi-agent debate.
These results suggest that naive debate protocols can inadvertently amplify errors rather than correcting them, underscoring the need for more principled approaches to multi-agent debate. Future debate frameworks should incorporate mechanisms that promote critical evaluation over consensus – such as encouraging agents to assess the soundness of others’ reasoning, integrating confidence estimates or credibility scores to weight contributions by expertise, and rewarding independent verification of claims. Training or incentive schemes might further discourage superficial agreement by penalizing unsupported conformity. Ultimately, improving the robustness of debate requires aligning models not merely to achieve consensus, but to engage in constructive epistemic disagreement – challenging peers when warranted and maintaining justified beliefs under social pressure.
Acknowledgments and Disclosure of Funding